### v1

# QUOTA: Quantifying Objects with Text-to-Image Models for Any Domain

Abstract
We tackle the problem of quantifying the number of objects by a generative text-to-image model. Rather than retraining such a model for each new image domain of interest, which leads to high computational costs and limited scalability, we are the first to consider this problem from a domain-agnostic perspective. We propose QUOTA, an optimization framework for text-to-image models that enables effective object quantification across unseen domains without retraining. It leverages a dual-loop meta-learning strategy to optimize a domain-invariant prompt. Further, by integrating prompt learning with learnable counting and domain tokens, our method captures stylistic variations and maintains accuracy, even for object classes not encountered during training. For evaluation, we adopt a new benchmark specifically designed for object quantification in domain generalization, enabling rigorous assessment of object quantification accuracy and adaptability across unseen domains in text-to-image generation. Extensive experiments demonstrate that QUOTA outperforms conventional models in both object quantification accuracy and semantic consistency, setting a new benchmark for efficient and scalable text-to-image generation for any domain.
1 Introduction
Text-to-image generative models [67, 40, 57, 7] have achieved remarkable success in generating high-quality images from textual descriptions. However, accurately controlling the count of objects in these generated images remains challenging, particularly when faced with text prompts that embody domains different from the ones seen during training. Recently, Zafar et al. [66] introduce prompt learning to refine the object counting prompts. While effective in controlled scenarios, this method struggles to maintain accurate object counts when confronted with varying domains. As a result, achieving domain-generalizable object quantification remains a crucial and unresolved challenge across various generative models, including diffusion and flow-matching models. Figure 1 illustrates this issue, showing how standard generative models like SDXL [41] struggle with accurate object quantification across different domains, whereas our approach provides consistent and accurate results across diverse domains.
Figure 1: Quantifying objects for any domain. The images outlined in red represent cases where the generated images have incorrect object counts. SDXL [41] struggles to control object counts across various domains, demonstrating inconsistencies even in domains typically encountered during training. IoCo by Zafar et al. [66], trained specifically on the photo domain, also faces challenges in generalizing object quantification to unseen styles. Our proposed QUOTA method achieves consistent and accurate object quantification for any domain.
Domain generalization in machine learning aims to create models that perform well on unseen domains, reducing the need for retraining and enabling greater adaptability across diverse settings. A wide range of approaches has been developed for this purpose, including methods that focus on learning domain-invariant features and leveraging diverse training environments [31, 9, 63]. Among these approaches, meta-learning has emerged as a promising technique, enabling models to adapt to new tasks by utilizing prior knowledge from related tasks [20, 60, 12]. Meta-learning has demonstrated success in improving domain generalization across various applications by optimizing models to adapt to different scenarios. Despite these advancements, there has been no exploration of meta-learning for domain-generalizable object quantification in text-to-image generation. In this paper, we fill this gap by proposing a novel meta-learning framework specifically designed to enable robust object quantification across diverse and unseen domains within text-to-image models.
Specifically, we make the following three contributions in this paper. First, we introduce a domain-agnostic setting for object quantification in generative text-to-image models, emphasizing the need for precise, domain-invariant object quantification across domains. This setting is essential for adapting models to diverse and unseen prompts without retraining, addressing both scalability and efficiency challenges. Second, we propose a dual-loop meta-learning strategy for prompt optimization, designed to make prompts domain-invariant. In the training phase, we divide text prompts into meta-train and meta-test domains. In the inner loop, we optimize prompt parameters on the meta-train domain to meet specific domain requirements. The outer loop refines these parameters on the meta-test domain, allowing the model to learn a prompt that generalizes across unseen domains. Third, we incorporate prompt learning with learnable quantification and domain tokens, enabling the model to capture stylistic variations effectively while preserving object quantification accuracy across domains. This approach allows for robust adaptation to new domains and unseen object classes without modifying the model architecture or retraining.
For evaluation, we adopt a new benchmark specifically designed for object quantification in domain generalization, enabling rigorous assessment of accuracy and adaptability across unseen domains in text-to-image generation.
Comparative evaluations against state-of-the-art methods demonstrate that QUOTA outperforms existing models in both object quantification accuracy and semantic alignment, setting a new benchmark for efficient, adaptable, and scalable text-to-image generation.
2 Related Work
The field of image generation has seen extensive study, from early advancements with GANs [15, 34, 42, 43, 59, 30, 68] to recent breakthroughs with diffusion models [19, 6, 45, 13, 24]. Diffusion models, in particular, have unlocked remarkable capabilities for generating high-quality, diverse images directly from natural language descriptions, exemplified by models like DALL-E 2 [46], Imagen [51], Parti [65], Make-A-Scene [13], Stable Diffusion (SD) [49], and CogView2 [8]. Our method builds on text-conditioned image generation, specifically exploring how to enhance object counting consistency across domains. Below, we review related areas in editing and controlling image generation models.
Personalized Image Generation. Personalization in image generation typically focuses on fine-tuning models to represent specific concepts from limited examples. Textual Inversion  [14], for instance, learns a new token to represent a concept using just 3-5 images, while DreamBooth [50] fine-tunes diffusion model parameters. However, adapting models to represent quantitative concepts, such as object counts, poses unique challenges as it involves precise numeric representation rather than qualitative identity. In contrast, our work focuses on achieving precise, domain-generalizable object quantification across varied styles and contexts without retraining, extending personalized image generation to handle quantitative control across different visual styles and settings.
Optimizing Prompt Correspondence. Recent studies have improved the alignment between text prompts and generated images. Prompt-to-Prompt [17] manipulates textual encoder attention for object replacement, while Null-inversion [36] allows similar edits on real images. Attend-and-Excite [5] ensures all objects in a prompt are depicted by optimizing attention maps. However, these methods generally do not address numeric accuracy in object counts, as prompt manipulation alone often fails to ensure precise quantities in diffusion outputs. Our work addresses this gap by introducing meta-optimized prompts that enable reliable control over object quantities across varied styles and contexts, establishing a foundation for robust object quantification in diverse visual settings.
Domain Generalization. Domain generalization (DG) aims to learn knowledge that transfers well to unseen target domains, allowing models to generalize effectively without access to target domain data during training. DG is typically studied under two main settings: multi-source domain generalization and single domain generalization (SDG). Multi-source DG utilizes multiple source domains to enhance robustness, while SDG—considered more realistic yet challenging—relies on a single source domain, making it harder due to limited data diversity. Early approaches to DG drew inspiration from domain adaptation methods, focusing on domain alignment to learn domain-invariant features. Techniques such as adversarial training [25, 53] and distribution distance minimization [32, 33] were employed to reduce feature bias across multiple source domains. Additionally, self-supervised learning [4, 26] and meta-learning [9, 55] have been explored to improve generalization in DG tasks. However, to the best of our knowledge, no prior work has specifically addressed domain-generalizable object quantification in text-to-image diffusion models. Our work introduces a novel setting and methodology to achieve domain-invariant object quantification across diverse and unseen domains. This new setting expands the scope of domain generalization research to the area of text-to-image generation, where accurate object quantification across varied domains is crucial.
Meta-Learning. Meta-learning, or learning to learn, seeks to enable rapid and efficient adaptation to new tasks by leveraging prior learning experiences, as reviewed by Hospedales et al. [21]. It has been applied across diverse areas, including designing adaptive loss functions [54, 2, 3], creating task-specific model initializations [12], and advancing few-shot learning techniques [28, 58, 56, 9, 10]. Meta-learning approaches are generally categorized into metric-based [56, 58, 61, 29], memory-based [35, 52, 37, 38, 20, 11], and gradient-based methods [48, 16, 39, 31]. Since the introduction of Model-Agnostic Meta-Learning (MAML) [12], gradient-based approaches have gained popularity, though they often encounter challenges with meta-overfitting due to limited meta-training tasks [1, 64, 69, 22, 23, 27]. Overall, meta-learning serves as a crucial foundation for our approach, enabling effective domain generalization in object quantification by facilitating adaptive prompt optimization across diverse text-to-image generation tasks.
3 Preliminaries
Text-Conditioned Diffusion Process. The diffusion process generates an image $x\sim p(x)$ from random noise through an iterative denoising procedure. Given an image $x_{0}\sim p(x)$, training involves creating noisy samples $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\cdot\epsilon_{t}$, where $\alpha_{t}$ are hyperparameters and $\epsilon_{t}\sim\mathcal{N}(0,\text{I})$. A neural network is trained to predict the added noise $\epsilon_{t}$ using the following objective:
$\mathbb{E}_{x,\epsilon_{t},t}\left[\|\hat{\epsilon}_{\theta}(x_{t},t)-\epsilon_{t}\|_{2}^{2}\right].$ (1)
For conditional denoising, a similar objective is applied, with each step conditioned on an external factor $y$:
$\mathbb{E}_{x,\epsilon_{t},t}\left[\|\hat{\epsilon}_{\theta}(x_{t},t,y)-\epsilon_{t}\|_{2}^{2}\right].$ (2)
In text-to-image diffusion, training involves reconstructing images conditioned on textual descriptions, which guide the denoising process to match the text. This requires a large dataset to generalize mappings between text and object quantities accurately. Some approaches modify noise predictions to align with specific object counts [66], although these often require counting models capable of handling noisy, intermediate data. In our work, we propose a method that adjusts the text conditioning embeddings using feedback from an off-the-shelf counting model.
Meta-Learning. Meta-learning aims to build models capable of quickly adapting to new tasks with limited data by utilizing prior experiences across various tasks [12]. Let $\mathcal{D}$ represent a meta-training dataset composed of training and validation sets across a range of tasks $\mathcal{T}$, i.e., $\mathcal{D}=\{\{D_{i}^{\text{tr}},D_{i}^{\text{val}}\}\}_{i\in\mathcal{T}}$, where $D_{i}^{\text{tr}}$ and $D_{i}^{\text{val}}$ are the training and validation datasets for the $i$-th task. Meta-learning can be formulated as a bi-level optimization problem:
$\min_{\phi}$ $\;\sum_{i\in\mathcal{T}}\mathcal{L}_{\text{val}}({\hat{\theta_{i}}}(\phi);D^{\text{val}}_{i})$ (3)
s.t. $\;{\hat{\theta_{i}}}(\phi)=\operatorname*{arg\,min}_{\theta_{i}}\mathcal{L}_{\text{train}}(\theta_{i};\phi,D^{\text{tr}}_{i}),\forall i\in\mathcal{T},$ (4)
where $\mathcal{L}_{\text{val}}$ and $\mathcal{L}_{\text{train}}$ represent the validation and training losses for the upper- and lower-level optimizations, respectively. In this setup, $\theta_{i}$ denotes the task-specific parameters for the $i$-th task, while $\phi$ are the meta-parameters shared across tasks. The lower-level optimization (Eq. (4)) adapts task-specific parameters $\theta_{i}$ for each task using the meta-parameters $\phi$ and the task-specific training set $D_{i}^{\text{tr}}$. Meanwhile, the upper-level optimization (Eq. (3)) aims to find meta-parameters $\phi$ that minimize the aggregated validation loss across tasks based on the optimized task-specific parameters $\hat{\theta_{i}}(\phi)$. Our work leverages this meta-learning framework to achieve domain-generalizable object quantification in text-to-image diffusion models, employing a dual-loop optimization approach to refine style-invariant prompts across diverse domains.
4 Methods
Figure 2: QUOTA’s dual-loop meta-optimization. The inner loop optimizes prompt parameters $e_{\text{count}}$ and $e_{\text{style}}$ across meta-train domains to adapt to domain-specific prompts, with losses computed using YOLOv9 for quantification accuracy and CLIP for semantic alignment. The outer loop refines these parameters on a meta-test domain, enhancing the model’s generalization to unseen domains. This dual-loop structure enables consistent object quantification and domain adaptation across diverse domains without the need for retraining.
In this section, we introduce quantifying objects with text-to-image models for any domain (QUOTA), our approach for quantifying object quantities across diverse domains in text-to-image generative models. QUOTA leverages a dual-loop meta-learning framework to create domain-invariant prompts that guide object quantification without retraining, making it adaptable to new and unseen domains. Through prompt learning, QUOTA optimizes domain-specific and quantity-specific parameters to accurately adjust object quantities based on textual prompts, even when encountering new stylistic variations. Our framework operates on images generated by existing text-to-image models, such as SDXL [41], using object quantities inferred by a pre-trained detection model as feedback for optimization.
4.1 Domain-generalizable object quantification
Accurately controlling the quantity of objects in text-to-image generative models is essential for producing realistic images, especially when adapting to diverse domains. To address the challenge of domain generalization, we introduce a novel training framework where multiple domains are leveraged to optimize the text-to-image generation process for consistent object quantification. Let $\mathcal{D}=\{\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}\}$ represent the set of domains available in the training phase, where $\mathcal{D}_{\text{train}}$ denotes the meta-train domains and $\mathcal{D}_{\text{test}}$ represents a held-out meta-test domain. During each iteration, we randomly select one domain from the training domains as the meta-test domain $\mathcal{D}_{\text{test}}$, while the remaining domains are assigned as the meta-train domains $\mathcal{D}_{\text{train}}$. This randomized selection allows the model to experience different combinations of training and testing domains, enhancing its robustness and adaptability to unseen domains. During training, images are generated based on text prompts in each domain using a pre-trained diffusion model, such as SDXL [41], without access to ground truth images. This setting enables us to simulate diverse training scenarios, preparing the model for domain shifts.
In the testing phase, our goal is to generalize to an unseen domain $\mathcal{D}_{\text{unseen}}$, where the model is expected to quantify objects accurately in images generated by the text-to-image model. To evaluate the object counts in these generated images, we employ two pre-trained models: YOLOv9 [62] and CLIP-S [44]. YOLOv9 serves as an object detector, providing a quantitative measure of the number of objects present in the image. Meanwhile, CLIP-S offers an alignment score between the generated image and the text prompt, helping to assess how well the generated image adheres to the prompt’s specifications, including object count. Notably, since the diffusion process does not inherently support direct control over object counts, we rely on a pre-trained object counting model that operates on completed images rather than intervening in the intermediate denoising steps. Thus, our method focuses on refining the domain and quantification parameters of the prompt, allowing it to adapt to the generated image characteristics across various domains and ensure accurate quantify object in the final output.
4.2 Dual-loop meta-optimization
QUOTA leverages a dual-loop meta-learning approach to achieve domain-invariant object quantification across diverse domains. This process involves an inner loop, which adapts the prompt parameters across multiple meta-train domains, and an outer loop, which refines these parameters on a single meta-test domain to encourage domain generalization. Each loop’s loss function incorporates two components: a quantification loss for object accuracy and a CLIP matching penalty to maintain semantic consistency. See Figure 2 for an overview.
Inner loop optimization. In the inner loop, given a set of meta-train domains $\mathcal{D}_{\text{train}}$, we optimize the prompt parameters $\theta$ across multiple domains to ensure the adaptability of the model. For each domain $d\in\mathcal{D}_{\text{train}}$, the composite loss function includes:
Quantification loss. Our inner loop optimization process utilizes a differentiable counting function, Count, which estimates the quantity of objects for a specified class in an image. The quantification loss for a target class $c$ is defined as the difference between the estimated and desired object counts, given by:
$\mathcal{L}_{\text{counting}}(x_{d},c,N)=\left\|\texttt{Count}(x_{d},c)-N\right\|,$ (5)
where $\texttt{Count}(x_{d},c)$ provides an estimate of the object count for class $c$ within the generated image $x_{d}$. This loss is calculated using a detection-based dynamic scale, leveraging YOLOv9 [62] to estimate the count of objects of class $c$ in the generated image $x_{d}$. The estimated count is given by:
$\text{Count}(x_{d},c)=\sum_{i=1}^{h\cdot w}\hat{\lambda}_{\text{scale},x_{d},\mathcal{D}}\cdot\Phi_{i}(x_{d},c),$ (6)
where $\hat{\lambda}_{\text{scale},x_{d}}$ is a dynamically adjusted scaling factor for differentiation continuity, and $\Phi_{i}(x_{d},c)$ represents the detection score for class $c$ at each pixel. This loss penalizes the difference between the target count $N$ and the detected count, ensuring accurate object quantification.
CLIP matching penalty. To ensure the image semantics remain intact, we add a CLIP-based penalty term:
$\mathcal{L}_{\text{semantic}}(x_{d},c;\theta)=\text{CLIP}(x_{d},\text{``A }s_{d}\text{ of }c\text{''}),$ ” ) , (7)
where $s_{d}$ is a domain descriptor specific to the domain $d$ (e.g., “photo,” “sketch,” “cartoon,” or “painting”), and CLIP provides a matching score between the generated image $x_{d}$ and the textual description “A [style] of [class],” reinforcing both semantic and stylistic alignment.
The inner loop loss $\mathcal{L}_{\text{inner}}$ is then computed by summing these losses across all meta-train domains:
$\mathcal{L}_{\text{inner}}(\theta)=\sum_{d\in\mathcal{D}_{\text{train}}}\mathcal{L}_{\text{counting}}(x_{d},c,N;\theta)+\lambda\mathcal{L}_{\text{semantic}}(x_{d},c;\theta),$ (8)
where $\lambda$ is a hyperparameter balancing the importance of object quantification accuracy and semantic consistency. This inner loop optimization updates the prompt parameters $\theta$ to adapt to the specific characteristics of the meta-train domains.
Outer loop optimization. The outer loop further refines the prompt parameters by evaluating them on the meta-test domain $\mathcal{D}_{\text{test}}$ to promote generalization to unseen domains. After optimizing the prompt parameters $\theta$ in the inner loop across meta-train domains, the outer loop computes the loss on the meta-test domain, defined as:
$\mathcal{L}_{\text{outer}}(\phi)=\mathcal{L}_{\text{counting}}(x_{\text{test}},c,N;\theta)+\lambda\mathcal{L}_{\text{semantic}}(x_{\text{test}},c;\theta),$ (9)
where $x_{\text{test}}$ is the generated image using the updated prompt parameters $\theta$ from the inner loop on the meta-test domain $\mathcal{D}_{\text{test}}$, and $\phi$ represents the global meta-parameters. This outer optimization step updates $\phi$, ensuring that the learned prompt parameters are robust enough to generalize to new, unseen domains, achieving accurate object quantification while maintaining semantic fidelity.
Through this dual-loop meta-optimization framework, QUOTA adapts the prompt parameters to handle domain shifts effectively, enabling consistent object quantification across various domains without compromising image quality or semantic alignment.
4.3 Learnable quantification and domain tokens
To achieve adaptive control over both object quantification and domain in text-to-image generation, we introduce prompt learning with two learnable tokens, $e_{\text{count}}$ and $e_{\text{style}}$. These tokens enable the model to flexibly adjust to different domains while maintaining accurate object quantification. This approach enhances the model’s ability to generalize across diverse domains by encoding information specific to object quantification and stylistic features directly within the prompt. Notably, all other model parameters remain fixed, allowing the model to achieve domain adaptability without altering the underlying architecture.
The process starts with an initial prompt format: “A $s_{d}$ of $N$ $c$,” where $s_{d}$ represents the domain (e.g., “photo,” “sketch,” “cartoon”), $N$ specifies the target object count, and $c$ denotes the object class. To exercise control over the generated image, we concatenate two newly added pseudo token embeddings, $e_{\text{count}}$ and $e_{\text{style}}$, to the text embedding. These embeddings are randomly initialized and are iteratively updated during training to adaptively encode information about object quantification and domain.
Unlike optimizing existing tokens, which may corrupt tokens with multiple meanings or associations, the introduction of these pseudo tokens allows the model to learn domain-specific adjustments without impacting predefined tokens. Formally, we optimize the embeddings $e_{\text{count}}$ and $e_{\text{style}}$ for each image generated, where the prompt parameters $\theta$ in our meta-optimization framework refer to these learnable tokens. In the inner loop, $e_{\text{count}}$ and $e_{\text{style}}$ are trained across meta-train domains to capture domain-specific prompt characteristics, resulting in domain-specific embeddings. The outer loop then refines these embeddings by adjusting the meta-parameters to generalize across domains, using the meta-test domain to guide the final optimization. The outer loop’s updated meta-parameters represent our optimized, domain-generalizable prompt embeddings, ensuring consistency in object quantification and domain adaptation even for unseen domains.
4.4 Testing phase
During testing on a new domain, we simply incorporate the domain and quantification information from the test text prompt by concatenating it with our learned tokens, $e_{\text{style}}$ and $e_{\text{count}}$. For example, if the input text prompt is “a cartoon of 6 apples,” the prompt fed to SDXL [41] becomes “a cartoon $e_{\text{style}}$ of 6 $e_{\text{count}}$ apples.” This approach allows us to generate images accurately aligned with the desired domain and object count without needing to re-optimize $e_{\text{style}}$ and $e_{\text{count}}$, thus saving time and ensuring efficient adaptation to new domains.
Photo Painting Cartoon Sketch Average
YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
w/o Meta-optimization 12.75 16.75 73.2 10.40 13.78 74.0 11.49 13.28 73.1 11.33 12.92 71.9 11.49 14.18 73.1
w/ Meta-optimization 10.09 13.98 74.3 8.04 11.04 75.9 9.48 11.92 74.8 9.15 11.69 74.1 9.19 12.16 74.8
Table 1: Impact of dual-Loop meta-optimization on object quantification and semantic alignment across different domains on the FSC benchmark. Lower values for MAE and RMSE indicate better counting accuracy, while higher CLIP-S scores represent better alignment with prompt semantics. Meta-optimization shows improvements across all metrics, especially in challenging domains such as Sketch, demonstrating its effectiveness in enhancing both object quantification and domain generalization.
5 Experiments
5.1 Experimental setup
Photo Painting Cartoon Sketch Average
YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
SDXL [41] 14.77 20.26 73.2 12.27 17.17 73.5 14.10 17.97 72.5 11.92 16.51 71.4 13.26 17.98 72.7
w/ $e_{style}$ 14.41 19.26 73.3 11.31 15.28 73.7 12.73 15.07 72.9 9.91 13.25 72.7 12.28 16.15 73.2
w/ $e_{count}$ 14.12 19.60 73.9 10.20 14.44 74.1 10.66 13.17 73.7 8.75 11.89 73.3 10.94 14.76 73.8
Ours (w/ $e_{count}$ $e_{style}$) 10.09 13.98 74.3 8.04 11.04 75.9 9.48 11.92 74.8 9.15 11.69 74.1 9.19 12.16 74.8
Table 2: Impact of learnable tokens on object quantification and semantic alignment across different domains on the FSC benchmark. Adding learnable tokens $e_{\text{count}}$ and $e_{\text{style}}$ improves YOLO-based MAE, RMSE, and CLIP-S scores, with the best performance achieved when both tokens are used together.
Figure 3: Comparison of object quantification results across domains with different optimization settings. This figure illustrates object quantification performance across various styles for three methods: SDXL [41], using only the inner loop in our framework, and our full QUOTA framework. The examples demonstrate how SDXL and the inner-loop-only setting struggle to achieve accurate quantification across domains, while meta-optimization achieves precise and consistent object quantification.
Photo Painting Cartoon Sketch Average
YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$
MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
SDXL [41] 14.77 20.26 73.2 12.27 17.17 73.5 14.10 17.97 72.5 11.92 16.51 71.4 13.26 17.98 72.7
Only train domains 12.75 16.75 73.2 10.40 13.78 74.0 11.49 13.28 73.1 11.33 12.92 71.9 11.49 14.18 73.1
IoCo [66] 7.72 12.21 74.4 11.36 15.68 73.1 12.40 15.93 72.7 10.80 14.21 71.9 10.57 14.51 73.0
QUOTA 10.09 13.98 74.3 8.04 11.04 75.9 9.48 11.92 74.8 9.15 11.69 74.1 9.19 12.16 74.8
Only test domain 5.68 8.63 75.9 4.61 7.12 76.8 5.31 7.86 75.9 5.94 8.53 75.3 5.38 8.04 76.0
Train domains & Test domain 6.43 9.05 74.7 7.61 10.79 76.1 8.43 10.96 75.1 8.75 11.27 74.7 7.81 10.52 75.2
Table 3: Performance comparison with various baseline models on object quantification and semantic alignment across different domains on the FSC benchmark. The methods and numbers shown in gray represent the upper bound, as these models have been trained using the test domain. Note that IoCo [66] is specifically trained on the Photo, which explains its high performance on Photo. In contrast, our QUOTA achieves strong results across all domains without having been explicitly trained on Photo, highlighting its domain-generalizable capabilities.
Figure 4: Generalization to unseen classes. Comparison of object quantification and style consistency between stable diffusion and our QUOTA, on both trained and unseen (reused) classes. While SDXL [41] struggles with unseen classes, QUOTA accurately controls object quantification and maintains semantic alignment across both trained classes (e.g., oranges, birds) and unseen classes (e.g., strawberries, crows). This demonstrates the robustness of QUOTA in adapting to previously unseen classes and domains.
Benchmark. We developed QUANT-Bench to assess object quantification performance across various classes, quantities, and visual domains. The classes in QUANT-Bench are based on the FSC dataset [47], a widely used dataset for object quantification that includes 19 categories such as common objects, vehicles, and animals. For each class, we generated 25 examples for quantities ranging from one to 25, yielding a total of 475 samples. To evaluate the model’s ability to generalize across different domains, we created images in four distinct domains: Photo, Painting, Cartoon, and Sketch. Both training and testing categories are drawn from the FSC dataset. To test each domain, we used a leave-one-out setting, where three domains are used as meta-train domains and the remaining domain as the meta-test domain. This setup allows us to measure how well the model generalizes object quantification to unseen domains by training with textual descriptions from the other three domains.
Metrics. To evaluate quantification accuracy, we use the MAE (Mean Absolute Error) and RMSE (Root Mean Square Error) metrics based on YOLOv9 [62]. These metrics are calculated by comparing the detected object counts to the target count in each image, with MAE capturing the average absolute deviation and RMSE highlighting larger errors through squared deviations. For semantic alignment assessment, we use a normalized CLIP matching score, defined as $\operatorname{CLIP-S}(p,x_{d})=w\cdot\max(\cos(p,x_{d}),0)$, where $p$ is the prompt and $x_{d}$ is the generated image [18]. Lastly, we calculate the average performance across different domains to assess each model’s consistency and adaptability to diverse contexts.
Implementation Details. We conducted training and evaluation on a single NVIDIA A6000 GPU with 48GB of memory. Training each token requires approximately two minutes, totaling around 30 iterations, using the SDXL framework [41]. For image quality, we find that a single denoising step is sufficient. The optimized quantification token can be reused without additional optimization time. We set $\lambda{=}5$, a learning rate of 0.01 for optimization, and the clipcount scaling hyperparameter $\lambda_{\text{scale}}$ to 60 for a static scale. We will make our code available.
Baselines. We establish a baseline using SDXL [41] backbone. We also compare with IoCo [66], which learns the quantification token only during training and considers images in the Photo domain exclusively.
5.2 Results
Benefit of meta-optimization. Table 1 demonstrates the impact of meta-optimization on object quantification accuracy and semantic alignment across different image styles. On average, meta-optimization achieves substantial improvements in both MAE and RMSE, highlighting enhanced object quantification accuracy and adaptability. The benefits are particularly pronounced in challenging domains, such as Cartoon, where meta-optimization leads to improved precision and consistency. Additionally, meta-optimization improves semantic alignment, as indicated by higher CLIP-S scores, reflecting closer alignment with prompt semantics. Figure 3 provides qualitative examples, comparing outputs from stable diffusion, an inner-loop-only optimization, and our full QUOTA framework. While inner-loop-only optimization offers some gains, it still struggles with consistent counts. In contrast, our complete meta-optimization approach, which includes both inner and outer loops, enables precise and domain-invariant object quantification across all styles, even in difficult cases like Sketch. These results confirm that meta-optimization significantly enhances the model’s ability to generalize quantification and semantic consistency across diverse domains.
Benefit of learnable tokens. Table 2 demonstrates the impact of adding learnable tokens $e_{\text{count}}$ and $e_{\text{style}}$ on object quantification accuracy and semantic alignment across different domains. The standard stable diffusion model, without any learnable tokens, shows limited adaptability, achieving lower accuracy across all domains. Introducing the $e_{\text{style}}$ token alone provides moderate improvements, indicating that adapting to domain style contributes positively to performance. Adding only the $e_{\text{count}}$ token leads to further enhancements, emphasizing the importance of object count adaptation. When both learnable tokens $e_{\text{count}}$ and $e_{\text{style}}$ are combined, as in QUOTA, the model achieves the best performance across all metrics and domains. This setup consistently enhances both counting accuracy and domain adaptability, highlighting the complementary effects of using both tokens. These results confirm that the joint use of $e_{\text{count}}$ and $e_{\text{style}}$ significantly strengthens the model’s ability to generalize object quantification across diverse domains while maintaining robust semantic alignment.
Figure 5: Qualitative comparison of generated images across different domains and classes. This comparison highlights the superior quantification accuracy and domain consistency achieved by QUOTA across diverse domains. The images outlined in red represent cases where the generated images have incorrect object quantification. The box in the Sketch domain showcases a challenging failure case for QUOTA, emphasizing the difficulty of accurately quantifying objects in highly stylized and intricate settings.
Generalization on unseen classes. In this experiment, we demonstrate the ability of our model, trained on 19 predefined classes, to generalize effectively to unseen classes using the optimized prompts generated by QUOTA. As illustrated in Figure 4, we tested the model on new classes not encountered during training, such as strawberries and crows. This generalization capability arises from our dual-loop training approach, where both the inner and outer loops are exposed to different domains and classes. By alternating between meta-train and meta-test domains with varying classes in each loop, our framework mimics the testing scenario of unseen class predictions, preparing the model for diverse and previously unseen content. Our approach successfully adapts to both the unseen classes and their respective domains, accurately controlling the object quantification and preserving semantic alignment. These results highlight QUOTA’s robustness in generalizing object quantification and domain consistency to new classes, demonstrating its potential for broad applicability in diverse scenarios.
Comparison with different baselines. To provide a comprehensive understanding of QUOTA ’s performance, we include a qualitative comparison in Figure 5 alongside the quantitative results in Table 3. Figure 5 illustrates generated images across various domains and classes, with QUOTA demonstrating superior quantification accuracy and domain consistency. In contrast, SDXL [41] and IoCo [66] struggle with incorrect object quantification across different domains, as indicated by the red outlines on images with quantification errors. QUOTA achieves consistent object counts close to the target quantities in most cases, while the baseline models frequently misestimate counts, especially in complex domains. Notably, IoCo’s reliance on training within the photo domain explains its higher performance in this domain but highlights its limitations in generalizing to unseen styles. The highlighted box in the Sketch domain shows a failure case for QUOTA, emphasizing the challenges in accurately quantifying objects within highly stylized and intricate settings. Together, these visual and numerical comparisons underscore QUOTA ’s capability to set a new benchmark for object quantification accuracy and domain adaptability in text-to-image generation.
Limitations and Future Work. While our approach demonstrates robust performance across various domains, it has limitations in particularly challenging domains, where improvements are less pronounced. This may be due to the inherent difficulty of accurately quantifying objects in abstract or highly stylized representations, although our method still outperforms previous approaches in these cases. Additionally, the use of a dual-loop meta-learning strategy introduces a higher computational cost during training, as it requires more iterations for optimizing domain-invariant prompts. However, at test time, our approach remains faster than traditional adaptation-based methods, such as IoCo [66], which require additional adaptation steps.
For future work, we plan to explore alternative object detectors to improve object quantification accuracy in challenging domains. Additionally, our method could be extended to other text-to-image generation models and applied to different tasks, such as fine-grained attribute control and scene layout generation, broadening its applicability within generative modeling.
6 Conclusion
In this work, we presented QUOTA, a novel optimization framework for domain-agnostic object quantification in generative text-to-image models. Unlike existing methods that require retraining for each new domain, QUOTA enables accurate object quantification across diverse and unseen domains through a dual-loop meta-learning strategy that optimizes a domain-invariant prompt. By integrating learnable quantification and domain tokens within prompt learning, our approach captures stylistic variations and maintains accuracy, even for object classes not seen during training. We introduced a new benchmark specifically for evaluating domain-generalizable object quantification, allowing rigorous assessment of model adaptability and quantification accuracy across domains. Extensive experiments demonstrate that QUOTA outperforms conventional models in both object quantification accuracy and semantic consistency, setting a new standard for efficient, scalable, and flexible text-to-image generation across any domain.
7 Acknowledgment
This work is financially supported by the Inception Institute of Artificial Intelligence, the University of Amsterdam, the Cisco gift funding, and the allowance Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy.

### v2

@@
 Further, by integrating prompt learning with learnable counting and domain tokens, our method captures stylistic variations and maintains accuracy, even for object classes not encountered during training.
-For evaluation, we adopt a new benchmark specifically designed for object quantification in domain generalization, enabling rigorous assessment of object quantification accuracy and adaptability across unseen domains in text-to-image generation.
+For evaluation, we adopt a new benchmark QUANT-Bench specifically designed for object quantification in domain generalization, enabling rigorous assessment of object quantification accuracy and adaptability across unseen domains in text-to-image generation.
 Extensive experiments demonstrate that QUOTA outperforms conventional models in both object quantification accuracy and semantic consistency, setting a new benchmark for efficient and scalable text-to-image generation for any domain.
@@
 Text-to-image generative models [37, 22, 31, 4] have achieved remarkable success in generating high-quality images from textual descriptions.
-However, accurately controlling the count of objects in these generated images remains challenging, particularly when faced with text prompts that embody domains different from the ones seen during training.
+However, accurately controlling the count of objects in these generated images remains challenging.
 Recently, Zafar et al.
-[66] introduce prompt learning to refine the object counting prompts.
+[35] introduced prompt learning to refine the object counting prompts.
 While effective in controlled scenarios, this method struggles to maintain accurate object counts when confronted with varying domains.
+Similarly, CountGen [1] improves counting accuracy by guiding the denoising process through internal object representations; however, it is restricted to a single domain and depends on additional modules for layout prediction.
 As a result, achieving domain-generalizable object quantification remains a crucial and unresolved challenge across various generative models, including diffusion and flow-matching models.
-Figure 1 illustrates this issue, showing how standard generative models like SDXL [41] struggle with accurate object quantification across different domains, whereas our approach provides consistent and accurate results across diverse domains.
-Figure 1: Quantifying objects for any domain.
-The images outlined in red represent cases where the generated images have incorrect object counts.
-SDXL [41] struggles to control object counts across various domains, demonstrating inconsistencies even in domains typically encountered during training.
-IoCo by Zafar et al.
-[66], trained specifically on the photo domain, also faces challenges in generalizing object quantification to unseen styles.
-Our proposed QUOTA method achieves consistent and accurate object quantification for any domain.
-Domain generalization in machine learning aims to create models that perform well on unseen domains, reducing the need for retraining and enabling greater adaptability across diverse settings.
+Figure 1 illustrates this problem, showing how standard generative models like SDXL [23] struggle with accurate object quantification across different domains, whereas our approach provides consistent and accurate results for any domain.
+Figure 1: Quantifying objects across domains.
+In each column, one domain is held out as the unseen test domain (shown in the column title), while the remaining three domains are used for training.
+For example, in the first column Photo is the test domain and the models are trained on Sketch, Cartoon, and Painting.
+Red boxes indicate incorrect object counts.
+SDXL [23] and IoCo [35] struggle with consistent quantification across domains.
+In contrast, QUOTA accurately controls object counts across all domains, including unseen ones.
+Domain generalization in machine learning aims to create models that perform well on unseen domains, reducing the need for retraining and enabling greater adaptability across diverse settings [38].
 A wide range of approaches has been developed for this purpose, including methods that focus on learning domain-invariant features and leveraging diverse training environments [16, 5, 34].
 Among these approaches, meta-learning has emerged as a promising technique, enabling models to adapt to new tasks by utilizing prior knowledge from related tasks [13, 32, 6].
-Meta-learning has demonstrated success in improving domain generalization across various applications by optimizing models to adapt to different scenarios.
 Despite these advancements, there has been no exploration of meta-learning for domain-generalizable object quantification in text-to-image generation.
-In this paper, we fill this gap by proposing a novel meta-learning framework specifically designed to enable robust object quantification across diverse and unseen domains within text-to-image models.
+In this paper, we fill this gap by proposing a novel meta-learning framework designed to enable robust object quantification across diverse and unseen domains within text-to-image models.
 Specifically, we make the following three contributions in this paper.
-First, we introduce a domain-agnostic setting for object quantification in generative text-to-image models, emphasizing the need for precise, domain-invariant object quantification across domains.
+First, we introduce a domain-agnostic setting for object quantification in generative text-to-image models, emphasizing the need for precise, domain-invariant object quantification.
 This setting is essential for adapting models to diverse and unseen prompts without retraining, addressing both scalability and efficiency challenges.
 Second, we propose a dual-loop meta-learning strategy for prompt optimization, designed to make prompts domain-invariant.
-In the training phase, we divide text prompts into meta-train and meta-test domains.
+In the training phase, we divide text prompts into meta-train and meta-validation domains.
 In the inner loop, we optimize prompt parameters on the meta-train domain to meet specific domain requirements.
-The outer loop refines these parameters on the meta-test domain, allowing the model to learn a prompt that generalizes across unseen domains.
+The outer loop refines these parameters on the meta-validation domain, allowing the model to learn a prompt that generalizes across unseen domains.
+This design mimics real-world scenarios where models must handle diverse domain shifts without prior exposure, ensuring adaptability to novel distributions in text-to-image generation.
 Third, we incorporate prompt learning with learnable quantification and domain tokens, enabling the model to capture stylistic variations effectively while preserving object quantification accuracy across domains.
 This approach allows for robust adaptation to new domains and unseen object classes without modifying the model architecture or retraining.
-For evaluation, we adopt a new benchmark specifically designed for object quantification in domain generalization, enabling rigorous assessment of accuracy and adaptability across unseen domains in text-to-image generation.
-Comparative evaluations against state-of-the-art methods demonstrate that QUOTA outperforms existing models in both object quantification accuracy and semantic alignment, setting a new benchmark for efficient, adaptable, and scalable text-to-image generation.
+For evaluation, we adopt a new benchmark QUANT-Bench specifically designed for object quantification in domain generalization, enabling rigorous assessment of accuracy and adaptability across unseen domains in text-to-image generation.
+Comparative evaluations against state-of-the-art methods demonstrate that QUOTA outperforms existing models in both object quantification accuracy and semantic alignment, setting a new benchmark for domain-generalizable text-to-image generation, enabling accurate object quantification across any domain without retraining.
 2 Related Work
-The field of image generation has seen extensive study, from early advancements with GANs [15, 34, 42, 43, 59, 30, 68] to recent breakthroughs with diffusion models [19, 6, 45, 13, 24].
-Diffusion models, in particular, have unlocked remarkable capabilities for generating high-quality, diverse images directly from natural language descriptions, exemplified by models like DALL-E 2 [46], Imagen [51], Parti [65], Make-A-Scene [13], Stable Diffusion (SD) [49], and CogView2 [8].
-Our method builds on text-conditioned image generation, specifically exploring how to enhance object counting consistency across domains.
-Below, we review related areas in editing and controlling image generation models.
+Text-to-image generation.
+The field of image generation has experienced significant evolution, transitioning from initial Generative Adversarial Network (GAN)-based approaches [8] to the more recent diffusion-based models [12].
+Early GAN methodologies laid the groundwork for generating images from textual descriptions, but often faced challenges related to training stability and output quality [8].
+In contrast, diffusion models have emerged as a robust alternative, demonstrating enhanced performance in producing high-fidelity images conditioned on text prompts [12].
+Notably, models like DALL·E 2 [24] and Imagen [27] have showcased the capability to generate photorealistic images from textual inputs.
+Concurrently, flow matching models have been introduced, offering advantages in training stability and sample quality by aligning continuous normalizing flows with target distributions [19].
+Our approach builds upon these advancements in text-conditioned image generation, focusing on enhancing object counting consistency across diverse domains.
 Personalized image generation.
-Personalization in image generation typically focuses on fine-tuning models to represent specific concepts from limited examples.
-Textual Inversion  [14], for instance, learns a new token to represent a concept using just 3-5 images, while DreamBooth [50] fine-tunes diffusion model parameters.
-However, adapting models to represent quantitative concepts, such as object counts, poses unique challenges as it involves precise numeric representation rather than qualitative identity.
-In contrast, our work focuses on achieving precise, domain-generalizable object quantification across varied styles and contexts without retraining, extending personalized image generation to handle quantitative control across different visual styles and settings.
-Optimizing Prompt Correspondence.
-Recent studies have improved the alignment between text prompts and generated images.
-Prompt-to-Prompt [17] manipulates textual encoder attention for object replacement, while Null-inversion [36] allows similar edits on real images.
-Attend-and-Excite [5] ensures all objects in a prompt are depicted by optimizing attention maps.
-However, these methods generally do not address numeric accuracy in object counts, as prompt manipulation alone often fails to ensure precise quantities in diffusion outputs.
-Our work addresses this gap by introducing meta-optimized prompts that enable reliable control over object quantities across varied styles and contexts, establishing a foundation for robust object quantification in diverse visual settings.
+Personalized image generation aims to fine-tune models for specific concepts using limited examples.
+Textual Inversion [7] learns a new token with just a few images, while DreamBooth [26] fine-tunes diffusion models for personalized content.
+However, adapting models for quantitative concepts, such as object counting, remains challenging as it requires precise numerical control rather than qualitative representation.
+Recent works have improved prompt-image alignment through techniques like Prompt-to-Prompt [9] for object manipulation, Null-inversion [20] for real-image edits, and Attend-and-Excite [3] to ensure prompt consistency.
+Our work introduces meta-optimized prompts that enable accurate, domain-generalizable object quantification without retraining, ensuring reliable control over object counts across diverse styles and contexts for consistent text-to-image generation.
+Object counting benchmarks.
+Several benchmarks have been proposed for evaluating object counting in text-to-image generation, including DrawBench[27], CoCoCount[1], and YOLO-Count [36].
+DrawBench evaluates multiple aspects of image generation, including cardinality, but does not address domain generalization.
+CoCoCount focuses on counting accuracy within specific datasets (e.g., COCO), but remains domain-specific and does not generalize across unseen domains.
+On the other hand, YOLO-Count introduces benchmarks like OpenImg7-New and Obj365-New, which assess a model’s ability to generalize to unseen categories.
+In contrast, our QUOTA benchmark focuses on precisely generating objects across unseen domains without retraining, providing a more scalable solution for real-world applications.
 Domain generalization.
-Domain generalization (DG) aims to learn knowledge that transfers well to unseen target domains, allowing models to generalize effectively without access to target domain data during training.
-DG is typically studied under two main settings: multi-source domain generalization and single domain generalization (SDG).
-Multi-source DG utilizes multiple source domains to enhance robustness, while SDG—considered more realistic yet challenging—relies on a single source domain, making it harder due to limited data diversity.
-Early approaches to DG drew inspiration from domain adaptation methods, focusing on domain alignment to learn domain-invariant features.
-Techniques such as adversarial training [25, 53] and distribution distance minimization [32, 33] were employed to reduce feature bias across multiple source domains.
-Additionally, self-supervised learning [4, 26] and meta-learning [9, 55] have been explored to improve generalization in DG tasks.
-However, to the best of our knowledge, no prior work has specifically addressed domain-generalizable object quantification in text-to-image diffusion models.
-Our work introduces a novel setting and methodology to achieve domain-invariant object quantification across diverse and unseen domains.
-This new setting expands the scope of domain generalization research to the area of text-to-image generation, where accurate object quantification across varied domains is crucial.
-Meta-Learning.
-Meta-learning, or learning to learn, seeks to enable rapid and efficient adaptation to new tasks by leveraging prior learning experiences, as reviewed by Hospedales et al.
-[21].
-It has been applied across diverse areas, including designing adaptive loss functions [54, 2, 3], creating task-specific model initializations [12], and advancing few-shot learning techniques [28, 58, 56, 9, 10].
-Meta-learning approaches are generally categorized into metric-based [56, 58, 61, 29], memory-based [35, 52, 37, 38, 20, 11], and gradient-based methods [48, 16, 39, 31].
-Since the introduction of Model-Agnostic Meta-Learning (MAML) [12], gradient-based approaches have gained popularity, though they often encounter challenges with meta-overfitting due to limited meta-training tasks [1, 64, 69, 22, 23, 27].
-Overall, meta-learning serves as a crucial foundation for our approach, enabling effective domain generalization in object quantification by facilitating adaptive prompt optimization across diverse text-to-image generation tasks.
+Domain generalization (DG) aims to enable models to generalize to unseen domains without access to target domain data during training.
+It is typically studied under two settings: multi-source DG, which leverages multiple source domains for robustness, and single-domain DG (SDG), which is more realistic but challenging due to limited data diversity.
+Early DG approaches focused on domain-invariant feature learning via adversarial training [14, 29], distribution alignment [17, 18], and self-supervised learning [2, 15].
+Meta-learning [5, 30] has also been explored to enhance generalization.
+However, no prior work has specifically addressed domain-generalizable object quantification in text-to-image diffusion models.
+We introduce a novel setting and methodology to achieve domain-invariant object quantification across diverse and unseen domains.
 3 Preliminaries
-Text-Conditioned Diffusion Process.
-The diffusion process generates an image $x\sim p(x)$ from random noise through an iterative denoising procedure.
-Given an image $x_{0}\sim p(x)$, training involves creating noisy samples $x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\cdot\epsilon_{t}$, where $\alpha_{t}$ are hyperparameters and $\epsilon_{t}\sim\mathcal{N}(0,\text{I})$.
-A neural network is trained to predict the added noise $\epsilon_{t}$ using the following objective:
-$\mathbb{E}_{x,\epsilon_{t},t}\left[\|\hat{\epsilon}_{\theta}(x_{t},t)-\epsilon_{t}\|_{2}^{2}\right].$ (1)
-For conditional denoising, a similar objective is applied, with each step conditioned on an external factor $y$:
-$\mathbb{E}_{x,\epsilon_{t},t}\left[\|\hat{\epsilon}_{\theta}(x_{t},t,y)-\epsilon_{t}\|_{2}^{2}\right].$ (2)
-In text-to-image diffusion, training involves reconstructing images conditioned on textual descriptions, which guide the denoising process to match the text.
-This requires a large dataset to generalize mappings between text and object quantities accurately.
-Some approaches modify noise predictions to align with specific object counts [66], although these often require counting models capable of handling noisy, intermediate data.
-In our work, we propose a method that adjusts the text conditioning embeddings using feedback from an off-the-shelf counting model.
-Meta-Learning.
-Meta-learning aims to build models capable of quickly adapting to new tasks with limited data by utilizing prior experiences across various tasks [12].
-Let $\mathcal{D}$ represent a meta-training dataset composed of training and validation sets across a range of tasks $\mathcal{T}$, i.e., $\mathcal{D}=\{\{D_{i}^{\text{tr}},D_{i}^{\text{val}}\}\}_{i\in\mathcal{T}}$, where $D_{i}^{\text{tr}}$ and $D_{i}^{\text{val}}$ are the training and validation datasets for the $i$-th task.
-Meta-learning can be formulated as a bi-level optimization problem:
-$\min_{\phi}$ $\;\sum_{i\in\mathcal{T}}\mathcal{L}_{\text{val}}({\hat{\theta_{i}}}(\phi);D^{\text{val}}_{i})$ (3)
-s.t. $\;{\hat{\theta_{i}}}(\phi)=\operatorname*{arg\,min}_{\theta_{i}}\mathcal{L}_{\text{train}}(\theta_{i};\phi,D^{\text{tr}}_{i}),\forall i\in\mathcal{T},$ (4)
-where $\mathcal{L}_{\text{val}}$ and $\mathcal{L}_{\text{train}}$ represent the validation and training losses for the upper- and lower-level optimizations, respectively.
-In this setup, $\theta_{i}$ denotes the task-specific parameters for the $i$-th task, while $\phi$ are the meta-parameters shared across tasks.
+Iterative object count optimization (IoCo).
+IoCo by Zafar et al.
+[35] introduces an iterative method to optimize object count accuracy in text-to-image models.
+Given a generated image $x$, an object category $c$, and a target object count $N$, their framework refines the generated output by minimizing the following objective:
+$\min_{\theta}\mathcal{L}_{\text{counting}}(x,c,N)+\lambda\mathcal{L}_{\text{semantic}}(x,c,N),$ (1)
+where the counting loss ensures the estimated count aligns with the target:
+$\mathcal{L}_{\text{counting}}(x,c,N)=\left|g_{\text{cnt}}(x,c)-N\right|,$ (2)
+with $g_{\text{cnt}}(x,c)$ representing the estimated number of objects in the generated image. To maintain the semantic integrity of the image, a CLIP-based regularization is used:
+$\mathcal{L}_{\text{semantic}}(x,c,N)=\text{CLIP}(x,\text{``A photo of }N\text{ }c\text{''}).$ (3)
+IoCo iteratively updates a learnable counting token $e_{\text{cnt}}$ to adjust the model’s internal representation of quantity:
+$e_{\text{cnt}}\leftarrow e_{\text{cnt}}+\alpha\frac{\nabla_{e_{\text{cnt}}}\mathcal{L}(x,c,N)}{\|\mathcal{L}(x,c,N)\|_{2}}.$ (4)
+Meta-learning for domain generalization.
+In this work, meta-learning aims to develop models capable of rapidly adapting to new tasks with limited data by leveraging knowledge from multiple tasks [6].
+In the context of domain generalization, meta-learning seeks to train models that generalize well to unseen domains by simulating domain shifts during training [16].
+Let $\mathcal{S}$ represent a training dataset comprising training and validation sets across various source domains $\mathcal{S}$, i.e., $\mathcal{S}=\{\mathcal{S}^{\text{tr}},\mathcal{S}^{\text{val}}\}$, where $\mathcal{S}^{\text{tr}}$ and $\mathcal{S}^{\text{val}}$ are disjoint subsets of the source domain set $\mathcal{S}$, with the former used for training and the latter for validation.
+The meta-learning process for domain generalization can be formulated as a bi-level optimization problem:
+$\min_{\phi}$ $\;\sum_{s^{\text{val}}\in\mathcal{S}^{\text{val}}}\mathcal{L}_{\text{val}}({\theta_{s^{\text{val}}}^{*}};s^{\text{val}}).$ (5)
+s.t.
+$\;{\theta_{s^{\text{val}}}^{*}}=\operatorname*{arg\,min}_{\theta_{s^{\text{tr}}}}\mathcal{L}_{\text{train}}({\theta_{s^{\text{tr}}}};s^{\text{tr}}),\quad s^{\text{tr}}\in\mathcal{S}^{\text{tr}},$ (6)
+where $\mathcal{L}_{\text{val}}$ and $\mathcal{L}_{\text{train}}$ denote the validation and training losses at the upper and lower levels, respectively.
+Here, $\theta_{s^{\text{tr}}}$} represents the domain-specific parameters for source domain $s^{\text{tr}}$, and $\phi$ denotes the shared meta-parameters.
 The lower-level optimization (Eq.
-(4)) adapts task-specific parameters $\theta_{i}$ for each task using the meta-parameters $\phi$ and the task-specific training set $D_{i}^{\text{tr}}$.
-Meanwhile, the upper-level optimization (Eq.
-(3)) aims to find meta-parameters $\phi$ that minimize the aggregated validation loss across tasks based on the optimized task-specific parameters $\hat{\theta_{i}}(\phi)$.
-Our work leverages this meta-learning framework to achieve domain-generalizable object quantification in text-to-image diffusion models, employing a dual-loop optimization approach to refine style-invariant prompts across diverse domains.
+(6)) adapts $\theta$ to each source domain using the training set $\mathcal{S}^{\text{tr}}$ and the meta-parameters $\phi$.
+The upper-level optimization (Eq.
+(5)) updates $\phi$ by minimizing the combined validation loss across source domains for the optimized domain-specific parameters $\theta_{s^{\text{val}}}^{*}$}.
+During testing, the trained model is evaluated on data from unseen target domains $\mathcal{T}$, which were not accessible during training.
+The goal is to assess the model’s ability to generalize to these novel domains without additional fine-tuning.
 4 Methods
-Figure 2: QUOTA’s dual-loop meta-optimization.
-The inner loop optimizes prompt parameters $e_{\text{count}}$ and $e_{\text{style}}$ across meta-train domains to adapt to domain-specific prompts, with losses computed using YOLOv9 for quantification accuracy and CLIP for semantic alignment.
-The outer loop refines these parameters on a meta-test domain, enhancing the model’s generalization to unseen domains.
-This dual-loop structure enables consistent object quantification and domain adaptation across diverse domains without the need for retraining.
-In this section, we introduce quantifying objects with text-to-image models for any domain (QUOTA), our approach for quantifying object quantities across diverse domains in text-to-image generative models.
+In this section, we introduce quantifying objects with text-to-image models for any domain (QUOTA).
 QUOTA leverages a dual-loop meta-learning framework to create domain-invariant prompts that guide object quantification without retraining, making it adaptable to new and unseen domains.
@@
 4.1 Domain-generalizable object quantification
-Accurately controlling the quantity of objects in text-to-image generative models is essential for producing realistic images, especially when adapting to diverse domains.
-To address the challenge of domain generalization, we introduce a novel training framework where multiple domains are leveraged to optimize the text-to-image generation process for consistent object quantification.
-Let $\mathcal{D}=\{\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}}\}$ represent the set of domains available in the training phase, where $\mathcal{D}_{\text{train}}$ denotes the meta-train domains and $\mathcal{D}_{\text{test}}$ represents a held-out meta-test domain.
-During each iteration, we randomly select one domain from the training domains as the meta-test domain $\mathcal{D}_{\text{test}}$, while the remaining domains are assigned as the meta-train domains $\mathcal{D}_{\text{train}}$.
-This randomized selection allows the model to experience different combinations of training and testing domains, enhancing its robustness and adaptability to unseen domains.
-During training, images are generated based on text prompts in each domain using a pre-trained diffusion model, such as SDXL [41], without access to ground truth images.
-This setting enables us to simulate diverse training scenarios, preparing the model for domain shifts.
-In the testing phase, our goal is to generalize to an unseen domain $\mathcal{D}_{\text{unseen}}$, where the model is expected to quantify objects accurately in images generated by the text-to-image model.
-To evaluate the object counts in these generated images, we employ two pre-trained models: YOLOv9 [62] and CLIP-S [44].
-YOLOv9 serves as an object detector, providing a quantitative measure of the number of objects present in the image.
-Meanwhile, CLIP-S offers an alignment score between the generated image and the text prompt, helping to assess how well the generated image adheres to the prompt’s specifications, including object count.
-Notably, since the diffusion process does not inherently support direct control over object counts, we rely on a pre-trained object counting model that operates on completed images rather than intervening in the intermediate denoising steps.
-Thus, our method focuses on refining the domain and quantification parameters of the prompt, allowing it to adapt to the generated image characteristics across various domains and ensure accurate quantify object in the final output.
+Accurately quantifying objects in text-to-image generative models is essential for ensuring alignment with specified numerical constraints.
+However, achieving consistent object counts across diverse domains is challenging due to the inherent variability in domain-specific styles and the lack of explicit numerical control in diffusion-based generation.
+Let $\mathcal{S}^{\text{tr}}$ represent a set of source domains used during training.
+Existing methods for object quantification in text-to-image diffusion models primarily focus on optimizing count accuracy within a fixed domain $s^{\text{tr}}\in\mathcal{S}^{\text{tr}}$ but overlook the challenge of generalization.
+These approaches assume that the model is trained and evaluated within the same domain, failing to account for domain shifts encountered when generating images in unseen distributions.
+As a result, a model trained on $\mathcal{S}^{\text{tr}}$ may overfit to domain-specific characteristics, leading to inconsistent quantification performance when applied to novel styles.
+In this work, we formally introduce the problem of domain-generalizable object quantification in text-to-image diffusion models.
+Our goal is to ensure that object quantification remains reliable when applied to unseen target domains $\mathcal{T}$, where object distributions, styles, and scene compositions may differ considerably from $\mathcal{S}^{\text{tr}}$.
+To the best of our knowledge, this is the first study that investigates how to optimize object counting mechanisms in a way that remains robust across unseen domains.
+Without explicitly addressing domain generalization, object quantification methods risk becoming unreliable when applied to new domains, limiting their practical utility in real-world scenarios.
+Figure 2: QUOTA’s dual-loop meta-optimization.
+The inner loop optimizes prompt parameters $e_{\text{count}}$ and $e_{\text{style}}$ across meta-train domains (Cartoon and Sketch) to adapt to domain-specific prompts, with losses computed using YOLOv9 for quantification accuracy and CLIP for semantic alignment.
+The outer loop refines these parameters on a meta-validation domain (Painting), enhancing the model’s generalization to unseen domains (Photo).
+This dual-loop structure enables consistent object quantification and domain adaptation across diverse domains without retraining.
 4.2 Dual-loop meta-optimization
-QUOTA leverages a dual-loop meta-learning approach to achieve domain-invariant object quantification across diverse domains.
-This process involves an inner loop, which adapts the prompt parameters across multiple meta-train domains, and an outer loop, which refines these parameters on a single meta-test domain to encourage domain generalization.
-Each loop’s loss function incorporates two components: a quantification loss for object accuracy and a CLIP matching penalty to maintain semantic consistency.
+QUOTA employs a dual-loop meta-optimization framework to achieve domain-invariant object quantification across diverse domains.
+This framework consists of an inner loop, which adapts the prompt parameters across multiple meta-train domains, and an outer loop, which refines these parameters on meta-validation domains to improve generalization.
+By dynamically sampling meta-train and meta-validation domains during training, the model learns to adapt to domain shifts and ensure robust object quantification.
 See Figure 2 for an overview.
+Meta-learning domain partitioning.
+We denote the meta-train and meta-validation domains as $S^{tr}$ and $S^{val}$, respectively.
+During training, we dynamically partition $\mathcal{S}$ into meta-train domains $\mathcal{S}^{\text{tr}}$ and meta-validation domains $\mathcal{S}^{\text{val}}$.
+The model first optimizes domain-specific prompt parameters $\theta_{s^{\text{tr}}}$} using $\mathcal{S}^{\text{tr}}$, refining object quantification within each domain.
+It is then validated on $\mathcal{S}^{\text{val}}$ to refine domain-invariant prompt parameters $\phi$, promoting generalization beyond seen distributions.
 Inner loop optimization.
-In the inner loop, given a set of meta-train domains $\mathcal{D}_{\text{train}}$, we optimize the prompt parameters $\theta$ across multiple domains to ensure the adaptability of the model.
-For each domain $d\in\mathcal{D}_{\text{train}}$, the composite loss function includes:
+In the inner loop, given a set of meta-train domains $\mathcal{S}^{\text{tr}}$, we optimize the prompt parameters $\theta$ across multiple domains to ensure the adaptability of the model.
+For each domain $s^{\text{tr}}\in\mathcal{S}^{\text{tr}}$, the composite loss function includes: quantification loss $\mathcal{L}_{\text{counting}}$ and CLIP matching penalty $\mathcal{L}_{\text{semantic}}$.
 Quantification loss.
-Our inner loop optimization process utilizes a differentiable counting function, Count, which estimates the quantity of objects for a specified class in an image.
+Our inner loop optimization uses a differentiable counting function, Count, which estimates the quantity of objects for a specified class in an image.
 The quantification loss for a target class $c$ is defined as the difference between the estimated and desired object counts, given by:
-$\mathcal{L}_{\text{counting}}(x_{d},c,N)=\left\|\texttt{Count}(x_{d},c)-N\right\|,$ (5)
-where $\texttt{Count}(x_{d},c)$ provides an estimate of the object count for class $c$ within the generated image $x_{d}$.
-This loss is calculated using a detection-based dynamic scale, leveraging YOLOv9 [62] to estimate the count of objects of class $c$ in the generated image $x_{d}$.
+$\mathcal{L}_{\text{counting}}(x_{s}^{\text{tr}},c,N)=\left\|g_{\text{cnt}}(x_{s}^{\text{tr}},c)-N\right\|,$ (7)
+where $g_{\text{cnt}}(x_{s}^{\text{tr}},c)$ provides an estimate of the object count for class $c$ within the generated image $x_{s}^{\text{tr}}$.
+This loss is calculated using a detection-based dynamic scale, leveraging YOLOv9 [33] to estimate the count of objects of class $c$ in the generated image $x_{s}^{\text{tr}}$.
 The estimated count is given by:
-$\text{Count}(x_{d},c)=\sum_{i=1}^{h\cdot w}\hat{\lambda}_{\text{scale},x_{d},\mathcal{D}}\cdot\Phi_{i}(x_{d},c),$ (6)
-where $\hat{\lambda}_{\text{scale},x_{d}}$ is a dynamically adjusted scaling factor for differentiation continuity, and $\Phi_{i}(x_{d},c)$ represents the detection score for class $c$ at each pixel.
+$\text{Count}(x_{s}^{\text{tr}},c)=\sum_{i=1}^{h\cdot w}\hat{\lambda}_{\text{scale},x_{s}^{\text{tr}},\mathcal{D}}\cdot\Phi_{i}(x_{s}^{\text{tr}},c),$ (8)
+where $\hat{\lambda}_{\text{scale},x_{s}^{\text{tr}},\mathcal{D}}$ is a dynamically adjusted scaling factor for differentiation continuity, and $\Phi_{i}(x_{s}^{\text{tr}},c)$ represents the detection score for class $c$ at the $i$-th bounding box [35].
 This loss penalizes the difference between the target count $N$ and the detected count, ensuring accurate object quantification.
-CLIP matching penalty. To ensure the image semantics remain intact, we add a CLIP-based penalty term:
-$\mathcal{L}_{\text{semantic}}(x_{d},c;\theta)=\text{CLIP}(x_{d},\text{``A }s_{d}\text{ of }c\text{''}),$ ” ) , (7)
-where $s_{d}$ is a domain descriptor specific to the domain $d$ (e.g., “photo,” “sketch,” “cartoon,” or “painting”), and CLIP provides a matching score between the generated image $x_{d}$ and the textual description “A [style] of [class],” reinforcing both semantic and stylistic alignment.
+CLIP matching penalty. To preserve image semantics, we add a CLIP-based penalty term:
+$\mathcal{L}_{\text{semantic}}(x_{s}^{\text{tr}},c,N)=\text{CLIP}(x_{s}^{\text{tr}},\text{``A }s_{d}\text{ of }N\text{ }c\text{''}),$ (9)
+where $s_{d}$ is a domain descriptor specific to the domain $s$ (e.g., “photo,” “sketch,” “cartoon,” or “painting”), and CLIP provides a matching score between the generated image $x_{s}^{\text{tr}}$ and the textual description “A [style] of N [class],” reinforcing both semantic and stylistic alignment.
 The inner loop loss $\mathcal{L}_{\text{inner}}$ is then computed by summing these losses across all meta-train domains:
-$\mathcal{L}_{\text{inner}}(\theta)=\sum_{d\in\mathcal{D}_{\text{train}}}\mathcal{L}_{\text{counting}}(x_{d},c,N;\theta)+\lambda\mathcal{L}_{\text{semantic}}(x_{d},c;\theta),$ (8)
+$\mathcal{L}_{\text{inner}}({\theta_{s^{\text{tr}}}})=\sum_{s^{\text{tr}}\in\mathcal{S}^{\text{tr}}}\mathcal{L}_{\text{counting}}(x_{s}^{\text{tr}},c,N;{\theta_{s^{\text{tr}}}})$ (10)
+$+\lambda\mathcal{L}_{\text{semantic}}(x_{s}^{\text{tr}},c,N;{\theta_{s^{\text{tr}}}}),$ (11)
 where $\lambda$ is a hyperparameter balancing the importance of object quantification accuracy and semantic consistency.
@@
 Outer loop optimization.
-The outer loop further refines the prompt parameters by evaluating them on the meta-test domain $\mathcal{D}_{\text{test}}$ to promote generalization to unseen domains.
-After optimizing the prompt parameters $\theta$ in the inner loop across meta-train domains, the outer loop computes the loss on the meta-test domain, defined as:
-$\mathcal{L}_{\text{outer}}(\phi)=\mathcal{L}_{\text{counting}}(x_{\text{test}},c,N;\theta)+\lambda\mathcal{L}_{\text{semantic}}(x_{\text{test}},c;\theta),$ (9)
-where $x_{\text{test}}$ is the generated image using the updated prompt parameters $\theta$ from the inner loop on the meta-test domain $\mathcal{D}_{\text{test}}$, and $\phi$ represents the global meta-parameters.
-This outer optimization step updates $\phi$, ensuring that the learned prompt parameters are robust enough to generalize to new, unseen domains, achieving accurate object quantification while maintaining semantic fidelity.
+The outer loop further refines the prompt parameters by evaluating them on the meta-validation domain $\mathcal{S}^{\text{val}}$ to promote generalization to unseen domains.
+After optimizing the prompt parameters $\theta_{s^{\text{val}}}^{*}$} in the inner loop across meta-train domains, the outer loop computes the loss on the meta-validation domain, defined as:
+$\mathcal{L}_{\text{outer}}(\phi)=\mathcal{L}_{\text{counting}}(x^{\text{val}}_{s},c,N;{\theta_{s^{\text{val}}}^{*}})$ (12)
+$+\lambda\mathcal{L}_{\text{semantic}}(x^{\text{val}}_{s},c,N;{\theta_{s^{\text{val}}}^{*}}),$ (13)
+where $x^{\text{val}}_{s}$ is the generated image using the updated prompt parameters $\theta_{s^{\text{val}}}^{*}$} from the inner loop, and $\phi$ denotes the shared meta-parameters, i.e., the global initialization of $e_{\text{count}}$ and $e_{\text{style}}$.
+The inner loop adapts these tokens to domain-specific requirements, while the outer loop updates their shared initialization to improve generalization across domains.
+Unlike the inner loop, where the text prompt follows the structure “A $s_{d}^{\text{tr}}$ of N cc”, the outer loop utilizes a different domain descriptor derived from $\mathcal{S}^{\text{val}}$, modifying the prompt format to “A $s_{d}^{\text{val}}$ of N cc”.
+This ensures that the model is exposed to different domain styles between training and validation, reinforcing robustness to domain shifts.
+By training on $\mathcal{S}^{\text{tr}}$ and validating on $\mathcal{S}^{\text{val}}$, the outer optimization step updates $\phi$, ensuring that the learned prompt parameters generalize effectively to new, unseen domains.
+This process achieves accurate object quantification while maintaining semantic fidelity.
 Through this dual-loop meta-optimization framework, QUOTA adapts the prompt parameters to handle domain shifts effectively, enabling consistent object quantification across various domains without compromising image quality or semantic alignment.
 4.3 Learnable quantification and domain tokens
-To achieve adaptive control over both object quantification and domain in text-to-image generation, we introduce prompt learning with two learnable tokens, $e_{\text{count}}$ and $e_{\text{style}}$.
-These tokens enable the model to flexibly adjust to different domains while maintaining accurate object quantification.
-This approach enhances the model’s ability to generalize across diverse domains by encoding information specific to object quantification and stylistic features directly within the prompt.
+To achieve adaptive control over both object quantification and the domain in text-to-image generation, we introduce prompt learning with two learnable tokens, $e_{\text{count}}$ and $e_{\text{style}}$.
+These tokens enhance the model’s ability to generalize across diverse domains by encoding information specific to object quantification and stylistic features directly within the prompt.
 Notably, all other model parameters remain fixed, allowing the model to achieve domain adaptability without altering the underlying architecture.
-The process starts with an initial prompt format: “A $s_{d}$ of $N$ $c$,” where $s_{d}$ represents the domain (e.g., “photo,” “sketch,” “cartoon”), $N$ specifies the target object count, and $c$ denotes the object class.
 To exercise control over the generated image, we concatenate two newly added pseudo token embeddings, $e_{\text{count}}$ and $e_{\text{style}}$, to the text embedding.
@@
 In the inner loop, $e_{\text{count}}$ and $e_{\text{style}}$ are trained across meta-train domains to capture domain-specific prompt characteristics, resulting in domain-specific embeddings.
-The outer loop then refines these embeddings by adjusting the meta-parameters to generalize across domains, using the meta-test domain to guide the final optimization.
+The outer loop then refines these embeddings by adjusting the meta-parameters to generalize across domains, using the meta-validation domain to guide the final optimization.
 The outer loop’s updated meta-parameters represent our optimized, domain-generalizable prompt embeddings, ensuring consistency in object quantification and domain adaptation even for unseen domains.
 4.4 Testing phase
-During testing on a new domain, we simply incorporate the domain and quantification information from the test text prompt by concatenating it with our learned tokens, $e_{\text{style}}$ and $e_{\text{count}}$.
+During testing on a new domain $\mathcal{T}$, we simply incorporate the domain and quantification information from the test text prompt by concatenating it with our learned tokens, $e_{\text{style}}$ and $e_{\text{count}}$.
 For example, if the input text prompt is “a cartoon of 6 apples,” the prompt fed to SDXL [23] becomes “a cartoon $e_{\text{style}}$ of 6 $e_{\text{count}}$ apples.” This approach allows us to generate images accurately aligned with the desired domain and object count without needing to re-optimize $e_{\text{style}}$ and $e_{\text{count}}$, thus saving time and ensuring efficient adaptation to new domains.
@@
 w/ Meta-optimization 10.09 13.98 74.3 8.04 11.04 75.9 9.48 11.92 74.8 9.15 11.69 74.1 9.19 12.16 74.8
-Table 1: Impact of dual-Loop meta-optimization on object quantification and semantic alignment across different domains on the FSC benchmark.
-Lower values for MAE and RMSE indicate better counting accuracy, while higher CLIP-S scores represent better alignment with prompt semantics.
+Table 1: Benefit of meta-optimization on object quantification and semantic alignment across different domains on QUANT-Bench.
 Meta-optimization shows improvements across all metrics, especially in challenging domains such as Sketch, demonstrating its effectiveness in enhancing both object quantification and domain generalization.
-5 Experiments
-5.1 Experimental setup
 Photo Painting Cartoon Sketch Average
@@
 SDXL [23] 14.77 20.26 73.2 12.27 17.17 73.5 14.10 17.97 72.5 11.92 16.51 71.4 13.26 17.98 72.7
-w/ $e_{style}$ 14.41 19.26 73.3 11.31 15.28 73.7 12.73 15.07 72.9 9.91 13.25 72.7 12.28 16.15 73.2
-w/ $e_{count}$ 14.12 19.60 73.9 10.20 14.44 74.1 10.66 13.17 73.7 8.75 11.89 73.3 10.94 14.76 73.8
-Ours (w/ $e_{count}$ $e_{style}$) 10.09 13.98 74.3 8.04 11.04 75.9 9.48 11.92 74.8 9.15 11.69 74.1 9.19 12.16 74.8
-Table 2: Impact of learnable tokens on object quantification and semantic alignment across different domains on the FSC benchmark.
-Adding learnable tokens $e_{\text{count}}$ and $e_{\text{style}}$ improves YOLO-based MAE, RMSE, and CLIP-S scores, with the best performance achieved when both tokens are used together.
-Figure 3: Comparison of object quantification results across domains with different optimization settings.
-This figure illustrates object quantification performance across various styles for three methods: SDXL [41], using only the inner loop in our framework, and our full QUOTA framework.
-The examples demonstrate how SDXL and the inner-loop-only setting struggle to achieve accurate quantification across domains, while meta-optimization achieves precise and consistent object quantification.
-Photo Painting Cartoon Sketch Average
-YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$
-MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
-SDXL [41] 14.77 20.26 73.2 12.27 17.17 73.5 14.10 17.97 72.5 11.92 16.51 71.4 13.26 17.98 72.7
-Only train domains 12.75 16.75 73.2 10.40 13.78 74.0 11.49 13.28 73.1 11.33 12.92 71.9 11.49 14.18 73.1
-IoCo [66] 7.72 12.21 74.4 11.36 15.68 73.1 12.40 15.93 72.7 10.80 14.21 71.9 10.57 14.51 73.0
-QUOTA 10.09 13.98 74.3 8.04 11.04 75.9 9.48 11.92 74.8 9.15 11.69 74.1 9.19 12.16 74.8
-Only test domain 5.68 8.63 75.9 4.61 7.12 76.8 5.31 7.86 75.9 5.94 8.53 75.3 5.38 8.04 76.0
-Train domains & Test domain 6.43 9.05 74.7 7.61 10.79 76.1 8.43 10.96 75.1 8.75 11.27 74.7 7.81 10.52 75.2
-Table 3: Performance comparison with various baseline models on object quantification and semantic alignment across different domains on the FSC benchmark.
-The methods and numbers shown in gray represent the upper bound, as these models have been trained using the test domain.
-Note that IoCo [66] is specifically trained on the Photo, which explains its high performance on Photo.
-In contrast, our QUOTA achieves strong results across all domains without having been explicitly trained on Photo, highlighting its domain-generalizable capabilities.
-Figure 4: Generalization to unseen classes.
-Comparison of object quantification and style consistency between stable diffusion and our QUOTA, on both trained and unseen (reused) classes.
-While SDXL [41] struggles with unseen classes, QUOTA accurately controls object quantification and maintains semantic alignment across both trained classes (e.g., oranges, birds) and unseen classes (e.g., strawberries, crows).
-This demonstrates the robustness of QUOTA in adapting to previously unseen classes and domains.
-Benchmark.
+w/ $e_{\text{style}}$ 14.41 19.26 73.3 11.31 15.28 73.7 12.73 15.07 72.9 9.91 13.25 72.7 12.28 16.15 73.2
+w/ $e_{\text{count}}$ 14.12 19.60 73.9 10.20 14.44 74.1 10.66 13.17 73.7 8.75 11.89 73.3 10.94 14.76 73.8
+Ours (w/ $e_{\text{count}}$ & $e_{\text{style}}$) 10.09 13.98 74.3 8.04 11.04 75.9 9.48 11.92 74.8 9.15 11.69 74.1 9.19 12.16 74.8
+Table 2: Benefit of learnable tokens on object quantification and semantic alignment across domains on QUANT-Bench.
+Adding learnable tokens $e_{\text{count}}$ and $e_{\text{style}}$ shows improvements across all metrics, with the best performance achieved when both tokens are used together.
+5 Experiments
+5.1 Experimental setup
+QUANT-Bench.
 We developed QUANT-Bench to assess object quantification performance across various classes, quantities, and visual domains.
-The classes in QUANT-Bench are based on the FSC dataset [47], a widely used dataset for object quantification that includes 19 categories such as common objects, vehicles, and animals.
+The classes in QUANT-Bench are based on the FSC dataset by Ranjan et al.
+[25], a widely used dataset for object counting that includes 19 classes such as common objects, vehicles, and animals.
 For each class, we generated 25 examples for quantities ranging from one to 25, yielding a total of 475 samples.
 To evaluate the model’s ability to generalize across different domains, we created images in four distinct domains: Photo, Painting, Cartoon, and Sketch.
-Both training and testing categories are drawn from the FSC dataset.
-To test each domain, we used a leave-one-out setting, where three domains are used as meta-train domains and the remaining domain as the meta-test domain.
+Both training and testing classes are drawn from the FSC dataset.
+To test each domain, we used a leave-one-domain-out setting, where two domains are used as meta-train domains for the inner loop, one domain is used as the meta-validation domain for the outer loop, and the remaining domain is held out as the unseen test domain.
 This setup allows us to measure how well the model generalizes object quantification to unseen domains by training with textual descriptions from the other three domains.
 Metrics.
-To evaluate quantification accuracy, we use the MAE (Mean Absolute Error) and RMSE (Root Mean Square Error) metrics based on YOLOv9 [62].
+To evaluate quantification accuracy, we use the Mean Absolute Error (MAE) and the Root Mean Squared Error (RMSE) based on YOLOv9 [33].
 These metrics are calculated by comparing the detected object counts to the target count in each image, with MAE capturing the average absolute deviation and RMSE highlighting larger errors through squared deviations.
-For semantic alignment assessment, we use a normalized CLIP matching score, defined as $\operatorname{CLIP-S}(p,x_{d})=w\cdot\max(\cos(p,x_{d}),0)$, where $p$ is the prompt and $x_{d}$ is the generated image [18].
-Lastly, we calculate the average performance across different domains to assess each model’s consistency and adaptability to diverse contexts.
+For semantic alignment assessment, we use the CLIP-S from Eq.
+(9), following [10].
+Conventional metrics such as FID [11] and Inception Score [28] are not appropriate for our QUANT-Bench.
+FID requires a real reference distribution, which is not available in our synthetic prompt space, and the Inception Score emphasizes sample diversity rather than alignment with the input prompt.
 Implementation details.
@@
 Training each token requires approximately two minutes, totaling around 30 iterations, using the SDXL framework [23].
-For image quality, we find that a single denoising step is sufficient.
+For image quality, we find a single denoising step suffices.
 The optimized quantification token can be reused without additional optimization time.
-We set $\lambda{=}5$, a learning rate of 0.01 for optimization, and the clipcount scaling hyperparameter $\lambda_{\text{scale}}$ to 60 for a static scale.
-We will make our code available.
-Baselines.
-We establish a baseline using SDXL [41] backbone.
-We also compare with IoCo [66], which learns the quantification token only during training and considers images in the Photo domain exclusively.
+We set $\lambda{=}5$, a learning rate of 0.01 for optimization, and the scaling hyperparameter $\lambda_{\text{scale}}$ to 60 for a static scale.
+Code will be released.
 5.2 Results
@@
 In contrast, our complete meta-optimization approach, which includes both inner and outer loops, enables precise and domain-invariant object quantification across all styles, even in difficult cases like Sketch.
-These results confirm that meta-optimization significantly enhances the model’s ability to generalize quantification and semantic consistency across diverse domains.
+These results confirm that meta-optimization considerably enhances the model’s ability to generalize quantification and semantic consistency across diverse domains.
 Benefit of learnable tokens.
 Table 2 demonstrates the impact of adding learnable tokens $e_{\text{count}}$ and $e_{\text{style}}$ on object quantification accuracy and semantic alignment across different domains.
-The standard stable diffusion model, without any learnable tokens, shows limited adaptability, achieving lower accuracy across all domains.
+The standard SDXL, without any learnable tokens, shows limited adaptability, achieving lower accuracy across all domains.
 Introducing the $e_{\text{style}}$ token alone provides moderate improvements, indicating that adapting to domain style contributes positively to performance.
@@
 When both learnable tokens $e_{\text{count}}$ and $e_{\text{style}}$ are combined, as in QUOTA, the model achieves the best performance across all metrics and domains.
-This setup consistently enhances both counting accuracy and domain adaptability, highlighting the complementary effects of using both tokens.
-These results confirm that the joint use of $e_{\text{count}}$ and $e_{\text{style}}$ significantly strengthens the model’s ability to generalize object quantification across diverse domains while maintaining robust semantic alignment.
-Figure 5: Qualitative comparison of generated images across different domains and classes.
-This comparison highlights the superior quantification accuracy and domain consistency achieved by QUOTA across diverse domains.
-The images outlined in red represent cases where the generated images have incorrect object quantification.
-The box in the Sketch domain showcases a challenging failure case for QUOTA, emphasizing the difficulty of accurately quantifying objects in highly stylized and intricate settings.
-Generalization on unseen classes.
-In this experiment, we demonstrate the ability of our model, trained on 19 predefined classes, to generalize effectively to unseen classes using the optimized prompts generated by QUOTA.
-As illustrated in Figure 4, we tested the model on new classes not encountered during training, such as strawberries and crows.
-This generalization capability arises from our dual-loop training approach, where both the inner and outer loops are exposed to different domains and classes.
-By alternating between meta-train and meta-test domains with varying classes in each loop, our framework mimics the testing scenario of unseen class predictions, preparing the model for diverse and previously unseen content.
-Our approach successfully adapts to both the unseen classes and their respective domains, accurately controlling the object quantification and preserving semantic alignment.
-These results highlight QUOTA’s robustness in generalizing object quantification and domain consistency to new classes, demonstrating its potential for broad applicability in diverse scenarios.
-Comparison with different baselines.
+These results confirm that the joint use of $e_{\text{count}}$ and $e_{\text{style}}$ strengthens the model’s ability to generalize object quantification across diverse domains while maintaining robust semantic alignment.
+Figure 3: Benefit of meta-optimization.
+We compare SDXL [23], our method with only the inner loop, and full QUOTA.
+Without meta-optimization, models struggle with accurate quantification.
+QUOTA achieves consistent counting across domains.
+Figure 4: Generalization to unseen classes.
+Comparison of SDXL [23] and QUOTA on both seen and unseen classes.
+QUOTA maintains accurate quantification and style consistency, while SDXL struggles with unseen classes, highlighting our method’s robustness.
+Photo Painting Cartoon Sketch Average
+YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$
+MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
+SDXL [23] 14.77 20.26 73.2 12.27 17.17 73.5 14.10 17.97 72.5 11.92 16.51 71.4 13.26 17.98 72.7
+Only train domains 12.75 16.75 73.2 10.40 13.78 74.0 11.49 13.28 73.1 11.33 12.92 71.9 11.49 14.18 73.1
+IoCo [35] 7.72 12.21 74.4 11.36 15.68 73.1 12.40 15.93 72.7 10.80 14.21 71.9 10.57 14.51 73.0
+QUOTA 10.09 13.98 74.3 8.04 11.04 75.9 9.48 11.92 74.8 9.15 11.69 74.1 9.19 12.16 74.8
+Target Domain 5.68 8.63 75.9 4.61 7.12 76.8 5.31 7.86 75.9 5.94 8.53 75.3 5.38 8.04 76.0
+Train & Target Domain 6.43 9.05 74.7 7.61 10.79 76.1 8.43 10.96 75.1 8.75 11.27 74.7 7.81 10.52 75.2
+Table 3: Comparison with others on object quantification and semantic alignment across domains in QUANT-Bench.
+Gray numbers denote target-domain upper bounds.
+While IoCo [35] is trained on Photo, QUOTA performs well across all domains without such supervision, highlighting its generalizability.
+Photo Painting Cartoon Sketch Average
+YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$ YOLO CLIP-S $\uparrow$
+MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$ MAE $\downarrow$ RMSE $\downarrow$
+SDXL [23] 14.07 19.80 72.8 20.64 27.05 73.0 11.71 15.40 72.1 18.39 24.03 71.1 16.20 21.57 72.3
+IoCo [35] 13.42 18.97 73.3 13.16 19.01 73.4 11.29 14.73 72.7 14.41 19.71 71.60 13.07 18.11 72.8
+QUOTA 13.72 19.20 73.1 10.99 15.24 74.5 9.94 12.22 73.2 12.05 14.68 72.3 11.68 15.33 73.3
+Table 4: Generalization to unseen classes.
+Comparison of SDXL [23], IoCo [35], and QUOTA on unseen classes across four target domains.
+QUOTA achieves consistently lower MAE and RMSE together with higher CLIP-S scores, demonstrating stronger quantification accuracy, semantic alignment, and robustness to domain shifts.
+Figure 5: Comparison with others by their generated images across different domains and object classes.
+This figure highlights the quantification accuracy and domain consistency of QUOTA.
+In the Painting domain (apples), QUOTA generates seven apples, but one appears significantly smaller, making it difficult for YOLOv9 to detect.
+Similarly, in the Sketch domain, the highly stylized and intricate setting challenges accurate object quantification, illustrating the limitations of detection-based counting in complex visual styles.
+Generalization to unseen classes.
+We further evaluate the ability of QUOTA, trained on 19 predefined classes, to generalize to classes not seen during training.
+Figure 4 provides qualitative examples, where we test on new categories such as strawberries and crows.
+While SDXL [23] struggles to maintain correct counts and style consistency, QUOTA successfully adapts to these unseen categories and domains, producing accurate quantification and faithful semantic alignment.
+Table 4 complements this analysis with quantitative results across four target domains.
+The unseen classes are grouped by semantic similarity to training classes (e.g., apples generalize to tomatoes, oranges, and strawberries; birds to crows, pigeons, and seagulls; sheep to zebras, horses, and cows).
+We report MAE, RMSE, and CLIP-S scores averaged over nine unseen classes with quantities from 1 to 25.
+Across all domains (Photo, Painting, Cartoon, and Sketch), QUOTA consistently outperforms SDXL [23] and IoCo [35], achieving substantially lower MAE and RMSE together with higher CLIP-S scores.
+These qualitative and quantitative results together highlight QUOTA’s robustness in extending object quantification to unseen classes while preserving semantic consistency, demonstrating its broad applicability across diverse domains and distributions.
+Methods MAE $\downarrow$ RMSE $\downarrow$ CLIP-S $\uparrow$
+Simple 9.19 12.16 74.8
+Complex 7.23 9.01 73.1
+Simple $\rightarrow$ complex 8.25 10.05 73.8
+Table 5: Effect of prompt complexity.
+Simple uses default prompts, while Complex includes GPT-4o-generated descriptions and trains the count token.
+Applying tokens from Simple to Complex (Simple $\rightarrow$ Complex) shows our method’s robustness to prompt variation.
+Effect of prompt complexity.
+To evaluate the impact of prompt complexity on object quantification, we compare three settings: Simple, Complex, and Simple $\rightarrow$ Complex, as shown in Table 5.
+The Simple setting uses default prompts, while Complex incorporates GPT-4o-generated [21] detailed descriptions and trains the count token as a learnable token, leading to improved performance.
+To assess the adaptability of our method, we apply the style and count tokens trained on simple prompts directly to complex prompts, which results in performance of the "Simple to Complex" setting.
+This demonstrates our method effectively generalizes across prompt complexities, maintaining robust object quantification without requiring additional retraining.
+Comparison with others.
 To provide a comprehensive understanding of QUOTA’s performance, we include a qualitative comparison in Figure 5 alongside the quantitative results in Table 3.
 Figure 5 illustrates generated images across various domains and classes, with QUOTA demonstrating superior quantification accuracy and domain consistency.
-In contrast, SDXL [41] and IoCo [66] struggle with incorrect object quantification across different domains, as indicated by the red outlines on images with quantification errors.
-QUOTA achieves consistent object counts close to the target quantities in most cases, while the baseline models frequently misestimate counts, especially in complex domains.
+In contrast, both SDXL [23] and IoCo [35] struggle with incorrect object quantification across different domains, as indicated by the red outlines on images with quantification errors.
+QUOTA achieves consistent object counts matching the target quantities in most cases, while the baseline models frequently misestimate counts, especially in complex domains.
 Notably, IoCo’s reliance on training within the photo domain explains its higher performance in this domain but highlights its limitations in generalizing to unseen styles.
-The highlighted box in the Sketch domain shows a failure case for QUOTA, emphasizing the challenges in accurately quantifying objects within highly stylized and intricate settings.
 Together, these visual and numerical comparisons underscore QUOTA’s capability to set a new benchmark for object quantification accuracy and domain adaptability in text-to-image generation.
-Limitations and Future Work.
-While our approach demonstrates robust performance across various domains, it has limitations in particularly challenging domains, where improvements are less pronounced.
-This may be due to the inherent difficulty of accurately quantifying objects in abstract or highly stylized representations, although our method still outperforms previous approaches in these cases.
-Additionally, the use of a dual-loop meta-learning strategy introduces a higher computational cost during training, as it requires more iterations for optimizing domain-invariant prompts.
-However, at test time, our approach remains faster than traditional adaptation-based methods, such as IoCo [66], which require additional adaptation steps.
-For future work, we plan to explore alternative object detectors to improve object quantification accuracy in challenging domains.
-Additionally, our method could be extended to other text-to-image generation models and applied to different tasks, such as fine-grained attribute control and scene layout generation, broadening its applicability within generative modeling.
 6 Conclusion
-In this work, we presented QUOTA, a novel optimization framework for domain-agnostic object quantification in generative text-to-image models.
-Unlike existing methods that require retraining for each new domain, QUOTA enables accurate object quantification across diverse and unseen domains through a dual-loop meta-learning strategy that optimizes a domain-invariant prompt.
-By integrating learnable quantification and domain tokens within prompt learning, our approach captures stylistic variations and maintains accuracy, even for object classes not seen during training.
-We introduced a new benchmark specifically for evaluating domain-generalizable object quantification, allowing rigorous assessment of model adaptability and quantification accuracy across domains.
-Extensive experiments demonstrate that QUOTA outperforms conventional models in both object quantification accuracy and semantic consistency, setting a new standard for efficient, scalable, and flexible text-to-image generation across any domain.
-7 Acknowledgment
-This work is financially supported by the Inception Institute of Artificial Intelligence, the University of Amsterdam, the Cisco gift funding, and the allowance Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy.
+We presented QUOTA, a novel optimization framework for domain-agnostic object quantification in text-to-image generation.
+Unlike existing methods that require retraining for each domain, QUOTA enables accurate quantification across diverse and unseen domains via a dual-loop meta-learning strategy that optimizes a domain-invariant prompt.
+By introducing learnable quantification and domain tokens into the prompt space, our approach effectively captures stylistic variations and maintains accuracy—even on object classes not seen during training.
+To facilitate systematic evaluation, we proposed a new benchmark tailored for domain-generalizable object quantification, enabling rigorous assessment of model adaptability and precision.
+Extensive experiments demonstrated that QUOTA outperforms standard baselines in both count accuracy and semantic alignment, establishing a new standard for flexible and scalable text-to-image generation.
+Limitations.
+Despite its strong cross-domain performance, QUOTA faces challenges in highly stylized or abstract domains where object boundaries are ambiguous.
+The dual-loop meta-learning introduces additional training overhead due to iterative prompt optimization, though test-time efficiency remains higher than adaptation-based methods such as IoCo [35].
+Furthermore, our reliance on external object detectors may introduce counting errors.
+Prior work [22] has shown that models like CLIP struggle with numerical reasoning, indicating that integrating counting-aware modules could further improve accuracy.
+Future work will explore stronger object detectors and extend QUOTA to other generative models, enabling applications in fine-grained attribute control and compositional scene generation.
+Disclaimer
+Funded by the European Union.
+Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission.
+Neither the European Union nor the European Commission can be held responsible for them.
+Acknowledgments
+This was supported by European Union’s Horizon Europe research and innovation programme under grant agreement number 101214398 (ELLIOT).
+Appendix A Performance Across Object Quantities
+Figure 6: MAE curves across object quantities for Painting and Cartoon domains. QUOTA significantly reduces error compared to SDXL [23] and IoCo [35], especially in the common range of 5–20 objects.
+Table 7 and Table 6 present the MAE results for 19 classes across two target domains: Painting and Cartoon, covering object quantities from 1 to 25.
+The results provide a detailed comparison between SDXL [23], IoCo [35], and QUOTA.
+In the Painting domain, QUOTA achieves the lowest MAE in 20 out of 25 quantities, consistently outperforming both IoCo [35] and SDXL [23].
+Similarly, in the Cartoon domain, QUOTA ranks first in 19 out of 25 cases, demonstrating its robustness in counting objects within stylized visual contexts.
+These findings highlight the strong performance of QUOTA in capturing complex visual variations and maintaining accurate object quantification across diverse classes and quantity ranges, further validating its effectiveness in real-world applications requiring precise object counting.
+Figure 6 complements these tables by visualizing the per-count MAE curves for the two target domains.
+The curves reveal that QUOTA’s largest gains concentrate in the practically most frequent range of 5–20 objects.
+In this interval, our method reduces the MAE by an average of 42.4% over SDXL [23] and 38.3% over IoCo [35] in the Painting domain, and by 36.0% and 28.8% respectively in the Cartoon domain.
+Below five objects, all three models already achieve very low errors, offering limited potential for further gains.
+In contrast, for counts above 20, the MAE remains high for all methods, likely due to increased visual complexity.
+These results indicate that QUOTA is particularly effective in the count range that matters most for real-world applications, while maintaining comparable performance on simpler cases.
+Appendix B Robust Generalization to Unseen Classes
+As shown in Figure 7, we train the models using the ’apples’ class in the painting target domain and evaluate their generalization on unseen classes with semantic similarity: ’tomatoes,’ ’oranges,’ and ’strawberries.’ While IoCo [35] performs well on the seen class during training, it fails to generalize effectively to unseen classes, leading to inaccurate quantification across all tested classes.
+Similarly, SDXL [23] struggles to adapt, producing inconsistent object counts and exhibiting misalignment with textual prompts.
+In contrast, our method, QUOTA, exhibits strong generalization, with accurate quantification across unseen classes.
+This highlights QUOTA ’s ability to adapt to unseen classes and preserve accuracy, even across stylistically distinct domains.
+These results highlight the effectiveness of QUOTA in addressing the challenges of cross-domain and cross-class generalization.
+Figure 7: Generalization to unseen classes. Comparison of object quantification between SDXL [23], IoCo [35], and our QUOTA, evaluated on both trained and unseen classes.